OpenStudy (anonymous):

Kevin and Randy Muise have a jar containing 81 coins, all of which are either quarters or nickles. The total value of the coins in the jar is $11.05. How many of each type of coin do they have?

OpenStudy (mertsj):

Write the equation that says the number of quarters plus the number of nickels is 81

OpenStudy (mertsj):

What is it?

OpenStudy (anonymous):

I'm not sure exactly what it is supposed to look like set up as an equation, that is the biggest problem I am having. I'm awful at word problems.

OpenStudy (mertsj):

Could you write the equation that says 5 plus 3 is 8?

OpenStudy (anonymous):

q + n = 81 -->q = 81 - n .25q + .05n = 11.05 now sub 81 - n in for q in the 2nd equation. Doing this will allow you to solve for n. Once n is solved for, substitute that answer in for n and solve for q. Once you have q and n, sub them both into either of the original equations, and if they come out equal you have the right answers. If they don't equal, then a mistake was made. Do you have any questions ?